OpenStudy (anonymous):

What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (anonymous):

Please explain how you did it too! I' like to learn!

OpenStudy (anonymous):

-4q+8=-2-16 You have to pick a side for your "q" on. So... -4q+8=-2q-16 +4q +4q Gives you 8=2q-16 Now add 16 to both sides to get rid of the 16 on the side with your q's 8=2q-16 +16 +16 24=2q then divide 2 from both sides to get your q alone 24/2=q 12=q

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

You're Welcome!